The simple interest on a certain sum of money for 3 years at 12%p.a is 6750. What will the compound interest on that sum at the same rate for the same period ?

Answer by satyareddy22(84)

S.I = (PTR)/100
S.I=6750 ,T =3 YRS,R=12% ,P= ?
SUBSTITUTING VALUES IN ABOVE FORMULA WE GET P=18750
NOW A=P(1+R/100)^T
P=18750 ,R=12%,T=3 YRS
A=18750(1+12/100)^3
= 18750(28/25)^3
= 26342.4
NOW C.I=A-P
=26342.4 - 18750 =7592.4
C.I = 7592.4